OpenStudy (anonymous):

If c = –5, how does the graph of y = x 2 + 2c compare to the graph of y = x 2 + c?

OpenStudy (anonymous):

they look the same the only difference is there position. y=x^2-10 is 5 steps below than y=x^2-5